# FBQuant: FeedBack Quantization for Large Language Models

### v1

@@ 1 Introduction
Existing optimization methods of weight-only quantization for LLMs can be broadly classified into three categories, as illustrated in Fig. 2, Clamping, Rotation, and Sub-braching: (a) Clamping methods Lin et al. (2024b); Shao et al. (2023) refine the quantization scale by restricting the value range, improving the representation of smaller weights while sacrificing outliers. (b) Rotation-based approaches Liu et al. (2024b); Lin et al. (2024a) apply rotation-equivalent transformations to shift quantization challenges from weights (which are quantized) to activations (which remain in high precision). (c) Sub-branching Li et al. (2023a, 2024) methods introduce a parallel branch alongside the quantized main path to compensate for quantization errors.
@@ 4.2 Differentiability for Layer-wise Reconstruction
which enables optimizing $\mathbf{\Sigma}$ by gradient descent. In practice, we implement $\mathbf{\Sigma}$ using low-rank adapters Hu et al. (2021), such that $\mathbf{\Sigma}=\mathbf{B}\cdot\mathbf{A}$. The sub-branches are applied to the linear layers in LLMs, such as Query, Key, Value, and Out projections in attention blocks, as well as Down, Gate, and Up projections in feed-forward networks (FFNs). During the optimization process, calibration data is fed into the model and each layer is optimized using gradient descent to minimize the reconstruction loss in Eq. 14. More Details of layer-wise reconstruction by FBQuant is described in algorithm 1.
@@ 5.1 Experimental Setup
Following previous works Frantar et al. (2022); Lin et al. (2024b), we employ 128 samples with a sequence length of 2048 in the subset of WikiText2 Merity et al. (2016) traning data for calibration. The perplexity results are tested on the WikiText2 validation set. The zero-shot evaluation is conducted using the open-source toolkit, i.e., Language Model Evaluation Harness Gao et al. (2024), which has been utilized by other baselines in their respective studies. The evaluation datasets include Arc-Challenge Clark et al. (2018), Arc-Easy Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), PIQA Bisk et al. (2020), WinoGrande ai2 (2019), and BoolQ Wang et al. (2019). We report the averaged accuracy in the main results.
@@ 5.2 Results and Analysis
Figure 7: Token throughput of the Llama2-7B model using FP16, INT4-Sub, INT4 and INT4-FBQuant tested on RTX 3090 GPU. Experiments were conducted with a batch size of 1. The rank parameter for INT4-Sub and FBQuant was set to 128. The token throughput is mearused by prefilling 256 and decoding 64 tokens.
@@ 6 Conclusions
Ethical Statement
There are no ethical issues.
Appendix A Appendix
A.1 Detailed Ill-posed Optimization
@@
A.2 Detailed Zero-shot Accuracy Results
We give detailed zero-shot results in Tabs. 3, 4, 5, 6, 7 and 8. The averaged accuracy is highlighted in bold. We apply various quantization methods to different LLMs. Quantization methods inculde Round-To-Nearest (RTN), GPTQ Frantar et al. [2022], AWQ Lin et al. [2024b], OmniQuant Shao et al. [2023], CALDERA Saha et al. [2024], SVDQuant Li et al. [2024], and our FBQuant. Models include Llama2 Touvron et al. [2023], Llama3 Dubey et al. [2024], and Qwen2.5 Yang et al. [2024] families. The benchmarks include Arc-Challenge Clark et al. [2018], Arc-Easy Clark et al. [2018], HellaSwag Zellers et al. [2019], MMLU Hendrycks et al. [2021], PIQA Bisk et al. [2020], WinoGrande ai2 [2019], and BoolQ Wang et al. [2019]. We report the averaged accuracy in the main results.

### v2

@@ 1 Introduction
Existing optimization methods of weight-only quantization for LLMs can be broadly classified into three categories, as illustrated in Fig. 2, Clamping, Rotation, and Sub-branching: (a) Clamping methods Lin et al. (2024b); Shao et al. (2023) refine the quantization scale by restricting the value range, improving the representation of smaller weights while sacrificing outliers. (b) Rotation-based approaches Liu et al. (2024b); Lin et al. (2024a) apply rotation-equivalent transformations to shift quantization challenges from weights (which are quantized) to activations (which remain in high precision). (c) Sub-branching Li et al. (2023a, 2024) methods introduce a parallel branch alongside the quantized main path to compensate for quantization errors.
@@ 4.2 Differentiability for Layer-wise Reconstruction
which enables optimizing $\mathbf{\Sigma}$ by gradient descent. In practice, we implement $\mathbf{\Sigma}$ using low-rank adapters Hu et al. (2021), such that $\mathbf{\Sigma}=\mathbf{B}\cdot\mathbf{A}$. The sub-branches are applied to the linear layers in LLMs, such as Query, Key, Value, and Out projections in attention blocks, as well as Down, Gate, and Up projections in feed-forward networks (FFNs). During the optimization process, calibration data is fed into the model and each layer is optimized using gradient descent to minimize the reconstruction loss in Eq. 14. More details of layer-wise reconstruction by FBQuant are described in algorithm 1.
@@ 5.1 Experimental Setup
Following previous works Frantar et al. (2022); Lin et al. (2024b), we employ 128 samples with a sequence length of 2048 in the subset of WikiText2 Merity et al. (2016) training data for calibration. The perplexity results are tested on the WikiText2 validation set. The zero-shot evaluation is conducted using the open-source toolkit, i.e., Language Model Evaluation Harness Gao et al. (2024), which has been utilized by other baselines. The evaluation datasets include Arc-Challenge Clark et al. (2018), Arc-Easy Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), PIQA Bisk et al. (2020), WinoGrande Sakaguchi et al. (2019), and BoolQ Wang et al. (2019). We report the averaged accuracy.
@@ 5.2 Results and Analysis
Figure 7: Token throughput of the Llama2-7B model using FP16, INT4-Sub, INT4 and INT4-FBQuant tested on RTX 3090 GPU. Experiments were conducted with a batch size of 1. The rank parameter for INT4-Sub and FBQuant was set to 128. The token throughput is measured by prefilling 256 and decoding 64 tokens.
@@ 6 Conclusions
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under Grant No. 2022YFB4400900, in part by the Strategic Industries and Key Technologies Project of Jiangsu Province under Grant BE2023020-3 and the Basic Research Program of Jiangsu Province under Grant BK20243042.
Appendix A Detailed Ill-posed Optimization
@@
Appendix B Detailed Zero-shot Accuracy Results
We give detailed zero-shot results in Tabs. 3, 4, 5, 6, 7 and 8. The averaged accuracy is highlighted in bold.